5. In the rod that has axis of rotation O, parallel forces F1 = 20N are applied,

F2 = 10N and F3 = 20N. AC = 2m, CO = 1m, OB = 1m. Find the resultant moment of the forces acting on the rod. What is the condition of the rod?


1
Expert's answer
2021-03-17T08:34:01-0400

I assume that the forces are acting in the same direction, and located as shown in figure below.



Then the moment that rotates the rod clockwise is:


"M_1 = F_3\cdot 1m = 20N\cdot m"

The moment that rotates the rod counterclockwise is:


"M_2 = F_1\cdot 2m + F_2\cdot 1m = 50N\cdot m"

The total moment is:


"M = M_2-M_1 = 30N\cdot m"

and the rotation is counterclockwise.


Answer. 30 N*m, counterclockwise.